OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (1, 3) and (3, 7)? y = −2x + 1 y = −2x − 1 y = 2x + 1 y = 2x − 1

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ 3}})\quad &({\color{red}{ 3}}\quad ,&{\color{blue}{ 7}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (texaschic101):

lets find the slope by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (1,3)...x1 = 1 and y1 = 3 (3,7)...x2 = 3 and y2 = 7 now we sub slope(m) = (7 - 3) / (3 - 1) slope(m) = 4/2 slope(m) = 2 now that we know the slope, we will find the y intercept by using the formula : y = mx + b (the m stands for the slope and the b stands for the y intercept) slope(m) = 2 you can use either of your points ... (1,3)...x = 1 and y = 3 now we sub 3 = 2(1) + b (by the way, we are solving for b, the y intercept) 3 = 2 + b 3 - 2 = b 1 = b so your equation is : y = 2x + 1

OpenStudy (anonymous):

wow thank you so much. really appreciate it